A planoconvex lens has a maximum thickness of 6 cm.When placed on a horizontal table with the curved surface in contact with the table surface,the apparent depth of the bottommost point of the lens is found to be 4 cm.if the lens is inverted such that the plane face of the lens is in contact with the surface of the table,the apparent depth of the center of the plane face is found to be [17/4] cm. The radius of curvature of the lens is ans-34 cmplz give me detail answerrates assured


Grade:10

1 Answers

Nathang
5 years ago
n=real depth/apparent depth
=6/4
=1.5
now take n1=1.5, n2=1, u=-6, v=-17/4 , R=?
n2/v-n1/u=(n1-n2)/R
1/(-17/4)-1.5/-6=(1.5-1)/R
R=34cm